OpenStudy (anonymous):

Express 12g into radian

OpenStudy (wolf1728):

12 g ???

OpenStudy (wolf1728):

Would that be grads? 400 grads = 360 degrees A grad is (360 / 400) degrees. a grad = 9/10 of a degree 2 PI radians = 400 grads

OpenStudy (wolf1728):

One grad = PI/200 radians So 12 grads = 12*PI/200 = 0.1884955592 radians